OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form.

OpenStudy (anonymous):

I love it.

OpenStudy (anonymous):

Standard form of a linear equation is Ax + By = C Point at (-1, 5), m=2 Step 1. Write the equation in Slope-Intercept form. y=mx+b 5=2(-1)+b Step 2. Solve for b 5=-2+b 7=b Now you have y=2x+7 (Slope intercept form) Step 3. Convert Slope Intercept to Standard Form y=mx+b Ax+By=C y=2x+7 -2x+y=7

OpenStudy (anonymous):

I hope you understand the concept ^^